### v1

# Reinforce Lifelong Interaction Value of User-Author Pairs for Large-Scale Recommendation Systems

Abstract.
Recommendation systems (RS) help users find interested content and connect authors with their target audience. Most research in RS tends to focus either on predicting users’ immediate feedback (like click-through rate) accurately or improving users’ long-term engagement. However, they ignore the influence for authors and the lifelong interaction value (LIV) of user-author pairs, which is particularly crucial for improving the prosperity of social community in short-video platforms. Currently, reinforcement learning (RL) can optimize long-term benefits and has been widely applied in RS. In this paper, we introduce RL to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) based on each interaction of UA pairs. To address the long intervals between UA interactions and the large scale of the UA space, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) and introduce an Adjacent State Approximation (ASA) method to construct RL training samples. Additionally, we introduce Multi-Task Critic Learning (MTCL) to capture the progressive nature of UA interactions (click → follow → gift), where denser interaction signals are leveraged to compensate for the learning of sparse labels. Finally, an auxiliary supervised learning task is designed to enhance the convergence of the RLIV-UA model. In offline experiments and online A/B tests, the RLIV-UA model achieves both higher user satisfaction and higher platform profits than compared methods.
Recommendation System, Lifelong Interaction Value, Reinforcement Learning, Sparse Cross-Request Interaction Markov Decision Process, Multi-Task Critic Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Computing methodologies Reinforcement learning.
1. Introduction
The recommendation system (RS) aims to explore users’ interested content, while helping content authors reach potential target users to accumulate fans and obtain profits (Zangerle and Bauer, 2022; Jesse and Jannach, 2021; Afsar et al., 2022). By promoting repeated interactions from both users and authors, RS helps build a more active ecosystem, leading to increased platform traffic and commercial returns (Azaria et al., 2013; Dias et al., 2008; Lu et al., 2014; Ras et al., 2017; Jannach and Adomavicius, 2017; Jannach and Jugovac, 2019; Jannach and Bauer, 2020).
Figure 1. The main process of the proposed RLIV-UA model optimizing the LIV of UA pair based on their interactions.
In the current field of RS, some research focuses on improving the accuracy of user immediate feedback prediction (Linden et al., 2003; He et al., 2014; Covington et al., 2016; Zhou et al., 2018; Gu et al., 2020), i.e. click-through rate (CTR), at each request by deep neural network (DNN) model, as shown in Fig. 1. Other studies concentrate on optimizing long-term engagement from the user’s perspective (Zhang et al., 2022; Xue et al., 2023; Chen et al., 2021; Wang et al., 2022; Cai et al., 2023; Zheng et al., 2018; Zou et al., 2019; Wang, 2021). Specifically, they utilize reinforcement learning (RL) to dynamically optimize session-level long-term cumulative rewards (Sutton, 2018; Van Hasselt et al., 2016; Fujimoto and Gu, 2021). However, the aforementioned research ignores the recommendation impact on authors and the Lifelong Interaction Value (LIV) of user-author (UA) pairs. Therefore, it is impossible to characterize the progressive changes of the lifelong UA relationship, which is essential for improving the stickiness (retention) of UA pairs and the prosperity of social community in short-video platforms.
(a) The relationship between the count of UA pairs with ”deep” relationship and the total revenue.
(b) The conversion rate of UA pairs with ”deep” relationship.
Figure 2. The revenue relevance and conversion rate of the lifelong UA relationship in Kwai app. Note that all data is collected in the second half of 2024 from Kwai app and scaled between 0 and 1.
Furthermore, we observe that the lifelong UA interactions have strong connection with the ultimate platform revenue. As shown in Fig. 2(a), the count of UA pairs with both follow and frequent gift (”deep”) relationship is positively correlated with the total revenue value. On the other hand, as users continue to interact with the authors with ”deep” relationship, its conversion rate will increase, as shown in Fig. 2(b). Hence, it provides insights to improve the platform profits by modeling the LIV of UA pairs.
To the best of our knowledge, this paper is the first to Reinforce Lifelong Interaction Value of User-Author pairs (RLIV-UA) using RL, based on each UA’s interaction and its corresponding cumulative reward. As shown in Fig. 1, the RLIV-UA model dynamically optimizes the LIV of an UA pair to improve the stickiness of both user U and author A progressively.
However, directly applying RL to model the LIV of UA pairs presents several challenges: First, compared with existing request-based RL applications in RS, the UA interactions may happen in requests with long time span. In practice, industrial RSs involve massive numbers of users and authors, resulting in an enormous and sparse UA state space, which limits the storage of all interaction traces of UA pairs over a long time period, such as three month.
To solve the above problems, we propose a novel Sparse Cross-Request Interaction Markov Decision Process (SCRI-MDP) and introduce an Adjacent State Approximation (ASA) method to construct RL training samples. Moreover, we propose a Multi-Task Critic Learning (MTCL) (Liu et al., 2023b) architecture to model gradually stronger UA relationships, such as click → follow → gift. Finally, due to the sample sparsity of each UA pair and high variance of labels, an auxiliary supervised learning task is designed to improve the stability and convergence of the RLIV-UA model.
Overall, the main contributions of this paper are as follows:
  • A novel LIV model, i.e. RLIV-UA, is proposed to reinforce the lifelong interaction value of UA pairs for large-scale RS.
  • Based on SCRI-MDP, a novel ASA method is introduced to construct RL training samples.
  • An MTCL architecture is proposed to model not only trivial interactions like click and watch time, but also ”deep” interactions like follow and gift. To improve the convergence of the RLIV-UA model, an auxiliary supervised learning task is designed.
  • Results in offline and online experiments show that the RLIV-UA model can improve both user engagement and author benefits, thus improving platform profits.
  • The RLIV-UA has been successfully deployed on two short-video applications, i.e. Kuaishou and Kwai, which include over 400 million daily active users.
Figure 3. The overall framework of the proposed RLIV-UA model.
2. PROBLEM FORMULATION
Existing RL methods in RS often model user behaviors as infinite request-level markov decision process (MDP) (Sutton, 2018). Specifically, the time interval between adjacent states $\Delta=1$ always holds. However, under the UA interaction space, the time interval between the same UA pair’s adjacent interactions satisfies $\Delta\geq 1$. For a specific UA pair, the interactions are usually sparse due to the large scale of candidate recommended items.
Therefore, we define a novel sparse cross-request interaction MDP (SCRI-MDP) to model the LIV of UA pairs. Formally, it is represented by a tuple of five elements $<\mathcal{S},\mathcal{A},P,\mathcal{R},\gamma>$:
  • State space $\mathcal{S}$: The state $\mathbf{s}_{ua}\in\mathcal{S}$ includes user $u$’s static features (e.g., user ID, location, gender), item and author $a$’s static features (e.g., item ID, author ID) and dynamic features of the $ua$ pair (e.g. cumulative gift count, cumulative watch time).
  • Action space $\mathcal{A}$: The action space $\mathcal{A}$ is defined as the candidate items in the ranking stage, and typically contains a few hundred items. The corresponding action $a\in\mathcal{A}$ is a recommended item list.
  • State transition probability distribution $P$: It is denoted as $P(\mathbf{s}_{ua}^{\prime}|\mathbf{s}_{ua},a)$ determined by the environment, where $\mathbf{s}_{ua}^{\prime}$ represents the next state of the same $ua$ pair.
  • Reward function $\mathcal{R}$: To model the progressive changes of UA relationship, several functions are designed to model the long-term reward of different immediate feedback between $ua$. Let $r_{ua,c},r_{ua,w},r_{ua,f},r_{ua,g}$ be the reward function of click, watch time, follow and gift labels respectively. $C=\{c,w,f,g\}$ denotes the target label set and $n=4$ denotes the cardinality of $C$.
  • Discount factor $\gamma$: $\gamma\in[0,1]$ is used to trade off instant rewards and future rewards.
We introduce a multi-task critic learning (MTCL) architecture to model the LIV of UA pairs by capturing the progressive changes of lifelong UA relationship. There are $n$ critic networks $Q_{\phi_{k}},k=1,...,n$ with $\phi_{k}$ as their trainable parameters to optimize the corresponding cumulative rewards. The final objective function is as follows:
(1) $\max_{\phi_{1},\cdots,\phi_{n}}\sum_{l\in C}\mathbb{E}\left[\sum_{t=0}^{\infty}{r_{ua,\,l}^{t}}\right]$
Each critic network is designed to learn from the states of UA pairs and corresponding rewards. Specifically, at one request of user $u$, RS recommends author $a$’s item $I_{i}$ which is the $t$-th interaction of $ua$, then the LIV of $ua$ pair is defined as:
(2) $Q(\mathbf{s}^{t}_{ua},I_{i})=r^{t}_{ua}+\gamma\max_{j\in\mathcal{I}}Q(\mathbf{s}^{t+1}_{ua},I_{j})$
where $\mathcal{I}$ is the item set containing all items produced by author $a$.
3. Methodology
In this section, we propose the RLIV-UA model to represent the SCRI-MDP in the UA pair state space. Firstly, due to the long time span between adjacent states of UA pair in the SCRI-MDP, we propose the Adjacent State Approximation (ASA) method to construct RL training samples. Then, we introduce the detailed network architecture of multi-task LIV networks and the final online deployment of the RLIV-UA model. Note that the overall framework of the RLIV-UA model is shown in Fig. 3.
3.1. Adjacent State Approximation
Since adjacent UA interactions happen in requests with a long time span, and the UA state space is extremely large with sparse UA interactions, the LIV of UA pairs is modeled as the SCRI-MDP. On one hand, the next state $s^{\prime}$ in the RL training sample $(s,a,r,s^{\prime})$ is not available until the next interaction of the same UA pair happens. On the other hand, it is impossible to store the interaction traces for all UA pairs in the industrial RS. Therefore, we design that the state space only contains user or author static features and the dynamic features between them:
(3) $\mathbf{s}^{t}_{ua}=\{\mathbf{f}_{u},\mathbf{f}_{a},\mathbf{f}_{ua}^{t}\}$
Specifically, the dynamic features $\mathbf{f}_{ua}^{t}$ in the current state $\mathbf{s}^{t}_{ua}$ is defined as the counts of several kinds of interactions between $ua$. Note that $\mathbf{f}_{ua}^{t+1}$ can be approximated based on user’s current rewards $r^{t}_{ua}$ and current dynamic UA features $\mathbf{f}_{ua}^{t}$:
(4) $\mathbf{f}^{t+1}_{ua}=\begin{cases}\mathbf{f}_{ua}^{t}+1,&\text{if }r^{t}_{ua}>0\\ \mathbf{f}_{ua}^{t},&\text{otherwise }\end{cases}$
Since the SCRI-MDP only focuses on the interactions and state transitions between UA pair like $ua$, it ignores interactions between user $u$ and items of other authors. Under the above assumption of the SCRI-MDP, the next state $\hat{\mathbf{s}}^{t+1}_{ua}$ can be derived by the dynamic features $\mathbf{f}^{t+1}_{ua}$.
3.2. Multi-Task LIV Networks
In order to model the progressive changes of lifelong UA relationship, we elaborately design 4 LIV networks based on the MTCL architecture, including Click, Watch time, Follow and Gift LIV networks. In particular, denser interaction signals (e.g. click and watch time) are used to compensate for sparse labels (e.g. follow and gift). For simplicity, we take one task tower as an example to explain the network structure.
As shown in Fig. 3, the current $ua$ state $\left\{\mathbf{f}_{u},\mathbf{f}_{a},\mathbf{f}_{ua}^{t}\right\}$ is fed into a shared Embedding Layer to obtain corresponding hidden embeddings $\mathbf{h}_{u}$, $\mathbf{h}_{a}$ and $\mathbf{h}_{ua}^{t}$. Taking the vector $\mathbf{H}^{t}_{ua}=\mathrm{concat}(\mathbf{h}_{u},\mathbf{h}_{a},\mathbf{h}_{ua}^{t})$ as the network input, the $Q_{\phi}(\mathbf{H}^{t}_{ua})$ is denoted as the final LIV of $ua$ at the $t$-th ineraction. To mitigate the deviation problem of value overestimation (Van Hasselt et al., 2016), double value networks and two corresponding target networks are used to output the minimum value:
(5) $Q_{\phi}(\mathbf{H}^{t}_{ua})=min(Q_{\phi^{1}}(\mathbf{H}^{t}_{ua}),Q_{\phi^{2}}(\mathbf{H}^{t}_{ua}))$
Then the corresponding loss function of a LIV network is defined as follows:
(6) $\mathcal{L}(\phi)$ $=$ $\mathbb{E}_{(\mathbf{s}^{t}_{ua},r^{t}_{ua},\mathbf{s}^{t+1}_{ua})\in D}[(Q_{\phi}(\mathbf{H}^{t}_{ua})-y)^{2}]$
(7) $y$ $=$ $r^{t}_{ua}+\gamma\max_{j\in\mathcal{I}}Q_{\phi^{\prime}}^{\prime}(\mathbf{H}^{t+1}_{ua})$
where $D$ indicates the real-time collected sample buffer, $y$ indicates the target output value of the LIV network, and $Q_{\phi^{\prime}}^{\prime}$ represents the output value of target network with same structure as LIV network $Q_{\phi}$. Note that the network parameter $\phi^{\prime}$ of $Q_{\phi^{\prime}}^{\prime}$ is periodically copied from $\phi$ of $Q_{\phi}$.
To reinforce the learning of sparse gift label of UA pairs in the Gift tower, an assistance neural network with two full-connected layers is designed to predict whether user $u$ will gift author $u$ at this interaction. As shown in Fig. 3, the user and author static hidden embeddings are input to the assistance network. And the binary cross entropy loss of the assistance gift binary classification goal is added to total loss.
3.3. Auxiliary Supervised Learning Network
Previous work (Liu et al., 2024) finds that the target value $y$ is often dominated by the inaccurate output of the target network $Q_{\phi^{\prime}}^{\prime}(\mathbf{H}^{t+1}_{ua})$ in practice, due to the instability of critic learning in RL. This problem reduces the effectiveness of the real reward $r^{t}_{ua}$ in guiding the learning of the value network, since it becomes relatively too small to provide meaningful learning signals. Furthermore, the large scale and extreme sparsity of the UA state space make the RL model even more difficult to converge.
Therefore, we introduce an auxiliary supervised learning network to regulate the learning of each LIV network $Q_{\phi}$, preventing a potential divergence of the RL model. Specifically, each LIV network is divided into two parts as follows:
(8) $Q_{\phi}(\mathbf{H}^{t}_{ua}):=\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})+\gamma\times\hat{V}_{\phi}(\mathbf{H}^{t}_{ua})$
where $\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})$ represents the current predicted reward between $ua$ at the $t$-th interaction with $\eta$ as its trainable parameters, and $\hat{V}_{\phi}(\mathbf{H}^{t}_{ua})$ is the future cumulative reward value excluding the current reward.
As real $r^{t}_{ua}$ is available based on the $t$-th interaction between $ua$, the predicted reward $\hat{r}_{\eta}(\mathbf{H}^{t}_{ua})$ can be learned by supervised loss. Incorporating with the aforementioned clipped double Q-learning (Fujimoto et al., 2018) shown in Eq. 5, the auxiliary supervised learning network improves the convergence of the RLIV-UA model.
Hence, the general loss of a LIV network is defined as:
(9) $\mathcal{L}_{Q}=huber\_loss(\hat{r}_{\eta}(\mathbf{H}^{t}_{ua}),r^{t}_{ua})+\sum_{2}^{k=1}(huber\_loss(Q_{\phi_{k}}(\mathbf{H}^{t}_{ua}),y)\$
where the first term denotes the loss for the auxiliary supervised learning network, and the second term denotes the original critic learning loss.
Overall, for the whole multi-task LIV networks, the final loss function is defined as follows:
(10) $\mathcal{L}=\sum_{l\in C}\mathcal{L}_{Q}^{l}+\mathcal{L}_{A}^{g}$
where $\mathcal{L}_{A}^{g}$ is the assistance gift loss and $\mathcal{L}_{Q}^{l}$ is the loss function, defined in Eq. 9, for each label in the target label set $C$, respectively. In practice, LIV scores from different task towers can be selectively applied based on the actual optimization goal, such as improving long-term user retention or maximizing platform revenue.
3.4. Online Deployment
Figure 4. System architecture of the RLIV-UA model in real industrial recommendation scenario.
In real industrial RS, the online system architecture of RLIV-UA model is illustrated in Fig. 4, including the offline training process and the online serving process. In the offline training process, we first utilize the label joiner to merge the UA state features $s$, action $a$, and the corresponding immediate feedback $r$ at timestamp $t$. Then we leverage the ASA method to approximate the next state $s^{\prime}$ and obtain RL training sample $(s,a,r,s^{\prime})$. In the online serving process, the RLIV-UA model is deployed during the ranking stage of RS to influence the final ordering of candidate items, whose number $N$ is typically less than 300.
Specifically, at each request of a user, the RLIV-UA model outputs corresponding LIVs for N candidate items. The candidate items are then ranked based on their LIV scores. Finally, the top K candidate items are selected as the final item list and are exposed to the user by order. Subsequently, the user interacts with each exposure item to return the feedback signals to RS.
4. Experiments
The proposed RLIV-UA model is evaluated in an offline simulated recommendation scenario to verify its performance compared with state-of-the-art models and the effectiveness of its each part. It is also applied in two real industrial recommendation scenarios to verify its effectiveness in large-scale industrial RS through online A/B tests.
4.1. Experimental Setup
4.1.1. Dataset and Evaluation Metrics
The KuaiRand (Gao et al., 2022) is used as the offline experimental dataset to train an offline user simulator. It is a public recommendation dataset containing 27,285 users and 32,038,725 items obtained from Kuaishou app. Therefore, there are hundreds of millions of UA interactions. The trained user simulator is used as the offline environment to mimic the users’ interaction with RS. In details, once simulator receives the recommended item, it returns immediate feedback like click, view and comments, etc. Then it determines whether to send the next request based on a quitting mechanism similar to that in work (Zhang et al., 2024).
We evaluate the performance of compared methods in three aspects, including user satisfaction, author benefits and platform profits:
  • User Satisfaction
    • Session Length: The number of requests in one session of a user with RS, which directly reflects the user satisfaction of the platform.
    • Watch Time: The accumulated watching time of all items watched by a user in one session.
    • CTR: The average click rate of all items recommended to a user in one session.
  • Author Benefits
    • Diversity: Quantifies the variety of content types in recommendations and is highly related to author benefits.
    • New Fans: The total number of new followers accumulated by authors.
  • Platform Profits
    • UA Count: The count of user-author pairs with ”deep” relationship which is defined as whether user has followed author, whether user has given author the most gifts, and other conditions.
    • Weekly Gifted Users: The number of gifted user in a week and it indicates the revenue scale of users in the platform.
    • Total Revenue: The important and ultimate metric to evaluate the platform revenue profits.
    • App Usage Time: Average time users spend on the app.
    • Weekly Retention: The stickness of user in a week.
4.1.2. Compared Methods
There are five models used as baseline in offline experiment and four variations of our model implemented in ablation experiment:
  • RankingModel: The classic ranking model in RS. Specifically, the DeepFM (Guo et al., 2017) is applied in offline experiments.
  • CQL(Kumar et al., 2020): An offline RL model that introduces a conservative constraint in the Q function update to limit the overly optimistic predictions of actions outside the data distribution.
  • DQN(Van Hasselt et al., 2016): A widely-used RL model that applies DNN as function approximator to estimate the Q-value of each action.
  • TD3(Fujimoto and Gu, 2021): A classic RL model which uses twin critics to reduce the bias, delays the update of policy and smooths the synchrony of target networks.
  • RLUR(Cai et al., 2023): A RL model that aims to optimize the weights of each predicted user feedback when ranking items under the long-term rewards with designed heuristic rewards to overcome the latency and sparsity of long-term rewards.
  • RLIV-UA(w/o MT): The proposed RLIV-UA model without multi-task LIV netowrks.
  • RLIV-UA(w/o MT & SL): The proposed RLIV-UA model without multi-task LIV netowrks and auxiliary supervised learning task.
  • RLIV-UA(w/o AL & SL): The proposed RLIV-UA model without assistance gift label network and auxiliary supervised learning task.
  • RLIV-UA(w/o SL): The proposed RLIV-UA model without auxiliary supervised learning task.
4.1.3. Implementation Details
Notably, all the models adopt the same hyperparameters listed in Table 1 for fair comparison.
Table 1. Hyperparameters of the compared models.
Hyper-parameter Value
Optimizer Adam
$\gamma$ Discount factor 0.9
$\tau$ Target network update rate 0.005
Learning rate of critic 1e-3
Learning rate of actor 1e-4
Batch size 1024
Train epochs 250
Hidden layer dimensions [64, 64]
The dimension of embedding layer 32
Learning rate of embedding layer 1e-3
Training steps per epoch 1e4
Training Platform PyTorch
The assistance gift label network is not applied in offline experiments, because there is no gift signals in KuaiRand dataset.
Besides, the platform profits metrics are only used in online A/B experiments. All models are trained to convergence and their results are the averaged performance of the last 10 epochs.
Moreover, we use the follow LIV and gift LIV in Kwai and use the watch time LIV in Kuaishou and offline experiments.
4.2. Performance Comparison
Table 2. Overall performance of different models in offline recommendation scenario.
Models Session Length Watch Time CTR Diversity
RankingModel 2.0132 59.4812 0.5948 0.0629
CQL 2.2660 58.0753 0.5125 0.1727
DQN 4.0610 49.5454 0.2247 0.7477
TD3 4.7820 52.7157 0.2228 0.8500
RLUR 6.6810 151.0550 0.4519 0.7206
RLIV-UA 12.8860 377.4867 0.5015 0.8827
Table 3. Overall performance of variations of the proposed RLIV-UA model in offline recommendation scenario.
Models Session Length Watch Time CTR Diversity
RLIV-UA(w/o MT & SL) 7.2940 188.0499 0.5229 0.7994
RLIV-UA(w/o MT) 9.2800 248.1804 0.5340 0.8746
RLIV-UA 12.8860 377.4867 0.5015 0.8827
The overall performance of different models in offline experiment is shown in Table 2. The traditional RankingModel achieves the best performance in CTR since it can predict which item has the greatest probability to be clicked. However, it is not suitable for improving the long-term user engagement such as session length and watch time. The offline model CQL learned from the historical samples can achieve some diversity. Compared with traditional RankingModel and the offline RL model CQL, most RL-based models achieve better performance in long-term metric session length at the expense of immediate feedback like low CTR, resulting in similar watch time. By adding another Q network learning from heuristic rewards, the RLUR model can improve all the long-term metrics including session length and watch time. The proposed RLIV-UA model achieves the best performance in session length and watch time and achieves the third high value in CTR, which reflects that the RLIV-UA model can balance immediate feedback and long-term feedback to improve long-term user engagement by modeling the LIV of UA pairs. Moreover, the RLIV-UA model achieves the best performance in diversity which reflects that modeling the LIV of UA pairs can more accurately recommend items of different authors to target users, rather than blindly recommend different items.
Figure 5. The learning process of RLIV-UA(w/o MT & SL) and RLIV-UA(w/o MT) over 10 rounds of training where the shaded areas correspond to the standard deviations.
4.3. Ablation Study
The overall performance of RLIV-UA variations in offline recommendation scenario are shown in Table 3. Firstly, compared with the RLIV-UA(w/o MT & SL) variation, the RLIV-UA(w/o MT) achieves relatively high improvement in session length, watch time and diversity, which reflects that the auxiliary supervised learning task can help model learn the LIV of UA pairs more accurately.
Furthermore, the RLIV-UA achieves the best performance under both session length, watch time and diversity metrics, which indicates the effectiveness of the multi-task critic learning architecture.
As shown in Fig. 5, with the auxiliary supervised learning task, the variance of RLIV-UA(w/o MT) is much lower than that of RLIV-UA(w/o MT & SL) under watch time, session length and CTR metrics. It demonstrates the learning process of RLIV-UA(w/o MT) model is more stable, and the auxiliary supervised learning task is effective for enhancing the stability of model training.
4.4. Online A/B Experiments
Firstly, the proposed RLIV-UA model and its variants are deployed on Kwai live-stream feed with over 100 million users and 1 million authors to improve platform revenue from July to October 2024. Through online A/B experiments, the RLIV-UA(w/o AL & SL), RLIV-UA(w/o SL) and RLIV-UA models are successively evaluated under the platform profits metrics compared to RankingModel as baseline. As shown in Table 4, RLIV-UA variations successively perform better in both revenue metrics and new fans metric, which illustrates the practical effectiveness of all parts of RLIV-UA model.
Table 4. Results of the revenue experiment on Kwai live-stream feed.
Models UA Count Weekly Gifted Users Total Revenue New Fans
RankingModel - - - -
RLIV-UA(w/o AL & SL) +2.002% +2.245% +11.903% +2.743%
RLIV-UA(w/o SL) +3.579% +3.417% +23.8065% +6.104%
RLIV-UA +5.849% +5.370% +38.6445% +8.282%
Table 5. Results of the watch time experiment on Kuaishou short-video feed.
Models Watch Time App Usage Time Weekly Retention
RankingModel - - -
RLIV-UA +0.131% +0.083% +0.021%
Figure 6. The output lifelong gift values of different grades of gift amount among UA pairs.
Moreover, the output lifelong gift values corresponding to different grades of gift amount among UA pairs are shown in Fig. 6, illustrating the high positive correlation between the output lifelong gift value and actual revenue.
The proposed RLIV-UA model is also deployed on Kuaishou short-video feed to improve long-term user engagement from November to December 2024. The results are listed in Table 5, which shows that the RLIV-UA model can improve the app usage time by 0.083% and the weekly retention by 0.021%. It should be emphasized that a 0.02% improvement holds statistical significance in the system. It is proved that the RLIV-UA model can not only improve platform revenue but also improve users’ long-term retention.
5. Related work
5.1. RL-Based Recommendation Systems
(Sunehag et al., 2015) is the earliest work that tries to alternate multitask learning ranking model with RL model using DQN to learn the value of all items in the recommended list. Similarly, Chen et al. (Chen et al., 2019) employ a policy-gradient approach in RS and Zhao et al. (Zhao et al., 2018) develop an actor-critic approach for recommending a page of items. However, they are not applied in a real-world recommendation environment with large amount of users and items. Then, more research (Gauci et al., 2018; Zheng et al., 2018) aims to apply the RL model in reality as a substitute with simple network structure. In order to handle the huge number of candidate items, SlateQ (Ie et al., 2019) is proposed to decompose the value of item list into the sum of value of each item under some assumptions. Recently, some literatures (Deffayet et al., 2023; Liu et al., 2023a) use contrastive learning to overcome the curse of dimensionality whose model structures are relatively more complex.
5.2. Long-Term User Engagement in Recommendation Systems
In order to consider the long-term user engagement rather than user’s immediate feedback, some research has increasingly focused on the sequential patterns of user behavior by employing temporal models, such as hidden Markov models and recurrent neural networks (Rendle et al., 2010; Campos et al., 2014; He and McAuley, 2016; Sahoo et al., 2012; Tan et al., 2016; Wu et al., 2017). Besides, some research (Gauci et al., 2018; Shani et al., 2005; Taghipour et al., 2007) use RL to make a long-term planning. However, all the methods are too complex to be applied in practice. Zou et al. (Zou et al., 2019) propose a hierarchical LSTM based Q network to model the complex user behavior and design an S-network to simulate the environment avoiding the instability. Chen et al. (Chen et al., 2021) inspired by exploration research (Mnih, 2016; Bellemare et al., 2016) in RL use a series of exploration methods to improve user experience. Wang et al. (Wang et al., 2022) carefully design the reward function through data analysis to connect the long-term rewards with immediate feedback. While Xue et al. (Xue et al., 2023) propose a framework for learning preferences from user historical behavior sequences, specifically using preferences to automatically train a reward function in an end-to-end manner. Considering all the above methods’ action is to select an item list which may be not practical when the number of item and user is large, Cai et al. (Cai et al., 2023) aim to optimize the weights of each predicted user feedback when ranking items under the long-term rewards with designed heuristic rewards to overcome the latency and sparsity of long-term rewards.
6. Conclusion
In this paper, we propose a novel lifelong interaction value model for user-author pairs, i.e. RLIV-UA, based on RL. Firstly, the interactions of UA pairs via RS is modeled as a sparse cross-request interaction markov decision process. To solve the long time interval and large scale of UA’s interactons, an adjacent state approximation method is designed to build the RL training sample. Besides, to capture the progressive changes of lifelong UA relationship, a multi-task critic learning architecture is employed to utilize denser interaction signals to compensate for sparse labels. Moreover, an auxiliary supervised learning task is designed to improve the convergence of the RLIV-UA model in large-scale RS. Finally, in both offline environments and online A/B tests, the experiment results show that the proposed RLIV-UA model performs better under both user satisfaction metrics and author benefits metrics, resulting in higher platform profits, compared with other models.
7. GenAI Usage Disclosure
We guarantee that no GenAI tools were used in any stage of the research, nor in the writing.